OpenStudy (anonymous):

The temperature of a cup of hot tea varies according to Newton's Law of Cooling: dT dt equals negative k times the quantity T minus A, where T is the temperature of the tea, A is the room temperature, and k is a positive constant. If the water cools from 100°C to 80°C in 1 minute at a room temperature of 60°C, find the temperature, to the nearest degree Celsius of the coffee after 4 minutes.

OpenStudy (anonymous):

https://gyazo.com/8aa6a713d1c912933a9dcf9581c5483f

OpenStudy (anonymous):

@jim_thompson5910 you taught me this a while back. I used your method and got 58 it would be nice if you can check it

OpenStudy (anonymous):

work with the temperature differences find the exponential model that goes through the points \((0,40)\) and \((1,20)\)

OpenStudy (abbycross167):

@satellite73 could you finish helping me with my question?

OpenStudy (anonymous):

initially the difference is 40, after one minute it is 20

OpenStudy (anonymous):

@satellite73 I am not sure what you're trying to get at I followed http://openstudy.com/users/ephemera#/updates/56468ffde4b0c3aa4a128f95

jimthompson5910 (jim_thompson5910):

ok I'm looking it over now

jimthompson5910 (jim_thompson5910):

@Ephemera can you post your work so far?

ganeshie8 (ganeshie8):

@Ephemera is your answer 58 ? the room temperature is 60, how can the temperature of tea go below room temperature ?

OpenStudy (anonymous):

Site was rally glitchy for me. And I know what I was doing wrong @ganeshie8 I used the same exact formula when in that question they ask for time, here they ask for temperature. According to your reasoning (very good reasoning might I add) the answer can't be 58. Please allow me a couple minutes to try and do it properly.

OpenStudy (anonymous):

Ok so this is what I have: \[\left| T-A \right| = e^c e^-kt\] T = 100 A= 80 t= 1 e^c = 60?

OpenStudy (anonymous):

\[\left| 100-80 \right| = 60e ^{-k}\]

OpenStudy (anonymous):

is it good so far?

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

OpenStudy (anonymous):

I'd solve for k then use k to find T?

jimthompson5910 (jim_thompson5910):

dt/(T-A) = -kdt ln(|T-A|) = -kt + C |T - A| = e^(-kt + C) |T - A| = e^C*e^(-kt) |T - 60| = e^C*e^(-kt) |100 - 60| = e^C*e^(-k*0) 40 = e^C e^C = 40 I'm not sure how you got e^C = 60

OpenStudy (anonymous):

hmm

ganeshie8 (ganeshie8):

Remember, \(T\) is a function of time \(t\). Your function for temperature at any given time should be : \[T(t) = 60 + Be^{-kt}\]

OpenStudy (anonymous):

What does B represent?

ganeshie8 (ganeshie8):

B is a constant

jimthompson5910 (jim_thompson5910):

B = e^C

ganeshie8 (ganeshie8):

plugin (0, 100) in above equation and you may solve B

OpenStudy (anonymous):

Same as e^c, I see it now

OpenStudy (anonymous):

So we need to solve for k then T now, correct?

ganeshie8 (ganeshie8):

T is not a constant

ganeshie8 (ganeshie8):

T is the temperature at any given time

OpenStudy (anonymous):

Are they not asking for T(4)?

ganeshie8 (ganeshie8):

Yes, they are

OpenStudy (anonymous):

And I can't solve for that without k correct?

ganeshie8 (ganeshie8):

Yes

OpenStudy (anonymous):

That is what I meant.

ganeshie8 (ganeshie8):

find the constants first using the given initial conditions

OpenStudy (anonymous):

I would use (1,80) to solve for k?

ganeshie8 (ganeshie8):

Yes, try finding B first though

OpenStudy (anonymous):

I thought be was equal to 40

OpenStudy (anonymous):

B*

OpenStudy (anonymous):

Or would it be a different scenario for 80?

jimthompson5910 (jim_thompson5910):

yes, B = 40

ganeshie8 (ganeshie8):

B is a constant, it doesn't change

OpenStudy (anonymous):

\[k= -\ln(1/2)\]

jimthompson5910 (jim_thompson5910):

or equivalently, k = ln(2)

OpenStudy (anonymous):

Yes using log properties.

OpenStudy (anonymous):

I'm getting T(4) = 700

jimthompson5910 (jim_thompson5910):

way too high

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

\[T(4) = 60 + 40e^{4\ln2}\]

jimthompson5910 (jim_thompson5910):

you should find that \[\Large T(t) = 60+40e^{-\ln(2)*t}\]

jimthompson5910 (jim_thompson5910):

you lost a negative in your exponent

OpenStudy (anonymous):

why would there be a negative?

OpenStudy (anonymous):

-ln(1/2) ------> ln(1/2)^-1

jimthompson5910 (jim_thompson5910):

\[\Large T(t) = 60+40e^{-kt}\] \[\Large T(t) = 60+40e^{-\ln(2)t}\]

OpenStudy (anonymous):

Which is ln 2

OpenStudy (anonymous):

Ohhhhh

OpenStudy (anonymous):

The negative in the formula, got you

OpenStudy (anonymous):

62.5 or 63

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks for both your help

jimthompson5910 (jim_thompson5910):

np